OpenStudy (anonymous):

A plastic bin contains red, yellow, and green key chains. Out of 350 key chains, 10% percent are yellow. Of the remaining key chains, 3/5 are green. How many red key chains are inside the bin? 26 35 126 189

hero (hero):

Howd I get here?

OpenStudy (hennawho):

find out how many key chains are green and yellow then subtract it from 350 to find the amount of red->>> yellow= .1 x 350 green= 3/5 x 350 red = 350 - green - yellow

OpenStudy (hennawho):

wait no thats not right. this is what u have to do...

OpenStudy (hennawho):

yellow= .1 x 350 = 35 Since 35 are yellow, the rest, 315, are red are green out of 315, 3/5 are green so, green = 3/5 x 315 = 189 so the rest are red 350-189-35= 126!!!!!